OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

(x) + (x+1) + (x+2) = 267 3x + 3 = 267 3x = 267 - 3 x = 264/3 x = 88 so th three numbers r 88, 89 and 90